OpenStudy (anonymous):

What is the value of time, t, for the given values of distance, d, and rate, r? d = 425 m, r = 50 m/s A. 8.5 m B. 8.5 s C. 21,250 m D. 21,250 s

jimthompson5910 (jim_thompson5910):

distance = rate*time d = r*t d/r = r*t/r ... Divide both sides by r. d/r = t t = d/r

jimthompson5910 (jim_thompson5910):

plug the given distance and speed into t = d/r to find the time

OpenStudy (anonymous):

I got 8.5

jimthompson5910 (jim_thompson5910):

correct 8.5 seconds

jimthompson5910 (jim_thompson5910):

the units for time are in seconds and not meters

OpenStudy (anonymous):

can u do more

jimthompson5910 (jim_thompson5910):

sure, one more

OpenStudy (anonymous):

Maria's family drove 140 miles to her grandparents' house and averaged 56 mi/h on the way there. On the return trip, they averaged 50 mi/ hr. What was the total time Maria's family spent driving to and from her grandparents' house? A. 2.5 h B. 2.6 h C. 5.2 h D. 5.3 h

jimthompson5910 (jim_thompson5910):

"Maria's family drove 140 miles to her grandparents' house and averaged 56 mi/h on the way there" how long did they take to go from their house to their grandma's house? use t = d/r

OpenStudy (anonymous):

so its 2.5 A

jimthompson5910 (jim_thompson5910):

that's just for the first part of the trip

jimthompson5910 (jim_thompson5910):

how long did they take coming back?

OpenStudy (anonymous):

wat do I do tho

jimthompson5910 (jim_thompson5910):

they travel 140 mi to get there, then 140 mi to get back

jimthompson5910 (jim_thompson5910):

so the distance d is the same each time

jimthompson5910 (jim_thompson5910):

the value of r coming back is 50

jimthompson5910 (jim_thompson5910):

use t = d/r to figure out how long it took them to come back home

OpenStudy (anonymous):

d?

jimthompson5910 (jim_thompson5910):

d = 140 the whole time

jimthompson5910 (jim_thompson5910):

the distance from home to grandmas house doesn't change

jimthompson5910 (jim_thompson5910):

how long does it take to go from home to grandma's house? how long does it take to go from grandma's house to home?

OpenStudy (anonymous):

2.8 but It says total time and I got d 5.3

jimthompson5910 (jim_thompson5910):

how did you get 5.3 ?

OpenStudy (anonymous):

I got 5.3 from 140/56=2.5 and 140/50=2.8 add them together and I got 5.3

jimthompson5910 (jim_thompson5910):

very good